OpenStudy (anonymous):

Solve for x 3x-6=7x+18 Show me the steps Thanks bunches

OpenStudy (anonymous):

well what we do to one side we have to do to the other side

OpenStudy (anonymous):

we need to get either 3x or 7x by its self

OpenStudy (anonymous):

@magan

OpenStudy (invinceble):

@magan

OpenStudy (anonymous):

Simplifying 3x + -6 = 7x + 18 Reorder the terms: -6 + 3x = 7x + 18 Reorder the terms: -6 + 3x = 18 + 7x Solving -6 + 3x = 18 + 7x Add '-7x' to each side of the equation. -6 + 3x + -7x = 18 + 7x + -7x Combine like terms: 3x + -7x = -4x -6 + -4x = 18 + 7x + -7x Combine like terms: 7x + -7x = 0 -6 + -4x = 18 + 0 -6 + -4x = 18 Add '6' to each side of the equation. -6 + 6 + -4x = 18 + 6 Combine like terms: -6 + 6 = 0 0 + -4x = 18 + 6 -4x = 18 + 6 Combine like terms: 18 + 6 = 24 -4x = 24 Divide each side by '-4'. x = -6 Simplifying x = -6

OpenStudy (anonymous):

sorry guys was busy

OpenStudy (anonymous):

thx gigi_here

OpenStudy (anonymous):

now just kinda close the question of u know whaqt i mean unless u need more help im here